# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v1

@@ 2.2 Related Work
In the realm of IDS-based ML incorporating XAI, a noteworthy instance is evident in the work presented in [38], the authors proposed the use of bidirectional SRU (Simple Recurrent Unit) with skip connections as a method for detecting attacks within IoMT networks. This approach effectively mitigates issues related to vanishing gradients, enhancing both training time and performance in attack detection. They compared the results achieved using two recurrent model variants, LSTM and GRU, leveraging the ToN-IoT dataset [18]. Furthermore, they compared their findings with prior studies and discovered that their solution outperformed in terms of accuracy, precision, recall, and F-score. Additionally, they conducted an analysis of important features using an XAI (Explainable AI) technique called Local Interpretable Model-Agnostic Explanations (LIME).
Examining the reviewed solutions reveals a absence of an integrated approach leveraging the advantages of ML for intrusion detection, FL for privacy protection, and XAI for interpretation and explanation within the IoMT. The proposed solution in this article stands out as the initial effort to combine these diverse technologies for IoMT systems security.
Ref
Methods
Datasets
Learning approach
use of the XAI
Optimizing FL parameter
[34]
@@ Methods
differential privacy contractive deep autoencoder for data fusion and quantum deep neural network method for ID
WUSTL-EHMS [15] and ICU [17]
Cent
No
No
[35]
DNN
CICIDS2017 [24]
FL
No
No
[36]
NN
MIMIC [36]
FL
No
No
[37]
hierarchical long-term memory
ToN-IoT and NSL-KDD [13]
FL
No
Yes except for learning strategies
[38]
bidirectional SRU with skip connections and LIME for XAI
ToN-IoT
cent
Yes
Non
[12]
PSO for features selections and AdaBoost for ID
NSL-KDD [13]
cent
No
No
[14]
RFE for features selections then MLP for AD
WUSTL-EHMS [15], ECU-IoHT [16], ICU [17], and ToN-IoT
Cent
No
No
[19]
@@
sSAE for dimensiality reduction then SGRU for ID
AWID [20]
Cent
No
No
[22]
@@
multidimensional DL model for features selection based on CNN, CNN-LSTM, and bidirectional LSTM then fully connected layers for ID
KISTI, KDDCup-99 [23], CICIDS2017 [24], and WSN-DS [25], as well as UNSW-NB15 [26]
Cent
No
No
[27]
PSO for features selections then DNN for ID
WUSTL-EHMS [15]
Cent
No
No
[28]
@@
integrates the LRGU technique into the MIFS. for features selection then ML for ID
WUSTL-EHMS [15]
Cent
No
No
[29]
PCA then GWO to minimize features then DNN for AD
@@
Kaggle intrusion data samples
Cent
No
No
[30]
ensemble learning that include NB, DT and RF
ToN-IoT
fog-cloud
No
No
[31]
@@
Deep Hierarchical Stacked Neural Networks
UNSW-BOT-IoT and UNSW-NB15 [26]
multi-cloud
No
No
[33]
ANN, RF, KNN, and SVM
WUSTL-EHMS [15]
Cent
No
No
@@ 3 METHODOLOGY FOR PROPOSED IDS MODEL
In this section, the IoMT system under consideration for the proposed framework is explored, followed by an outline of the FL process, which involves local training at the end device level and aggregation processes at the server level. Subsequently, the employed XAI method is detailed. The objective of this framework is to ensure that the collection, transfer, and processing of patient data are conducted securely, respecting their privacy and aligning with international standards in the medical domain.
@@ 3.1 IoMT Systems
In layer 1, various types of medical devices equipped with sensors and actuators are employed to collect vital information and administer medications to patients that suffer from chronic disease and also it can be used for fall detection for elderly persons or measuring the performance of athletes. These devices can be categorized into four types [39, 40]: implanted devices within the body, wearable devices, ambient devices capturing environmental data, and stationary devices found within hospitals. Given the energy limitations, wireless connections are established between these medical devices and mobile devices using low, or ultra-low-power wireless communication technologies such as Bluetooth, Zigbee, or NFC, thereby overcoming communication constraints [41].
@@
Finally, in layer 3, the centralized medical server takes charge of handling messages transmitted from the mobile devices and relaying them back to the patients. The server must be equipped with high computational capacity to effectively handle incoming and outgoing communications as well as perform in-depth analysis of the received data using AI methods. Additionally, it features a cloud server for intelligent decision-making, aggregating and storing additional patient medical data. The collected data is accessible to doctors, patients, and the pharmacy department through an online interface or smartphone. Integration with Electronic Health Record (EHR) and Electronic Medical Record (EMR) systems ensures easy access to information and provides notifications for uploaded or received health data. The figure 2 provides an overview of the described system [46].
@@ 3.2 Local Training Process
In this context, the inclination is towards employing supervised learning in DL for IDS, leveraging ANN for real-time anomaly detection. This architecture, simpler and less intricate than alternative models, not only enables adaptability and maintainability but also delivers outstanding performance. Its capability to conduct real-time detection, as showcased in section 4, underscores its efficiency. Furthermore, the model’s straightforward design not only helps in curbing energy consumption within personal devices but also contributes to reducing communication costs within the FL environment, which occur due to exchanging model weights during the FL process instead of row data. There are three types of neural layer in an ANN: an input layer, one or more hidden layers and an output layer. Each neuron contains a threshold and connections to other neurons with weighted connections. Neurons get activated if their cumulative weight surpasses the threshold, transmitting signals to the subsequent layer [48].
The rectified linear unit introduced in [49] serves as the activation function for the hidden layer due to its use of the MAX function (1), enabling faster computation and helping prevent overfitting, thus enhancing performance.
@@
The He-Initializer presented in [50] was preferred as it works well with the ReLu activation function (2).
@@
For the output layer, the sigmoid activation function was utilized (3), which is suitable for binary classification tasks.
@@
Where $x$ represent the input to the sigmoid function. It can be the weighted sum of the inputs from the previous layer.
In compiling the model, cross-entropy was employed as the loss function (4) and Adam described in [51] as the optimizer (5).
@@
where $y_{i}$ is the i-th element of the true distribution, and $p_{i}$ is the i-th element of the predicted distribution.
@@
Where $g_{t}$ is the gradient, $\alpha$ is the learning rate, $\beta_{1}$ and $\beta_{2}$ are the moment parameters, and $\epsilon$ is a small constant to avoid division by zero.
Adam’s advantage lies in its low memory requirement, making it suitable for large-scale data or parameter-heavy problems. This method combines the strengths of two popular optimization techniques: adagrad, effective in handling sparse gradients, and RMSProp, known for its efficacy in dealing with non-stationary objectives [51].
@@ 3.3 FL Porcess
Improving the security of the IoMT is crucial, and FL emerges as a pivotal solution. FL facilitates model training by allowing the exchange of locally trained model weights among end-devices, avoiding the transmission of raw data. This innovative approach upholds data privacy standards, aligning seamlessly with regulations such as HIPAA, and its decentralized nature mitigates the vulnerability of a centralized structure.
@@
initialize model $w_{0}$ with he initializer
$R\leftarrow$ Number of Round of communication
$E\leftarrow$ Number of local epoch
$Fr\leftarrow$ Fraction fit
$M\leftarrow$ Number of Patient
$\eta\leftarrow$ learning rate
for $R=1,2,3\dots$ do
     $C\leftarrow$ Random Set of $M\times Fr$ clients
     for patient $k\subset C$ in parallel do
         $w_{(t+1)}^{k}\leftarrow$ Patient Client Update $(k,w_{t},E)$
@@
     $w_{(t+1)}\leftarrow\sum_{k+1}^{k}\frac{\eta_{k}}{\eta}w_{(t+1)}^{k}$
@@
Algorithm 2 Patient Client Update $(k,w_{t},E)$
@@
for $e=0$ to $e=E-1$ do
     $w\leftarrow$ use Adam optimizer to update $w$ to minimize the binary cross-entropy loss function
end forreturn $w$ to server
@@
In Step 1, the security administrator initializes the global model with random weights at the hospital server. Personal devices selected for FL download the initial weights to initialize their local models. In Step 2, medical sensors send captured data to paired personal devices using short-range wireless communication. In Step 3, the local model undergoes training to update weights, utilizing a specified number of epochs as outlined in Algorithm 2, once personal devices have accumulated a sufficient amount of data. This strategy is designed to minimize the number of communication rounds needed for model convergence, thereby reducing communication costs through the optimization of bandwidth usage. In Step 4, updated weights of local models are sent to the hospital server via HTTPS communication. Step 5 sees the hospital server aggregating these weights to update the global model by summing the received weights and dividing by the number of selected participants in a communication round, thereby creating the global model as depicated in formula (6). Step 6 involves the global server sending updated weights to personal devices, and Steps 2-6 are repeated until the weight modifications become insignificant or the model converges. Once the model achieves the desired performance in Step 7, it is deployed for intrusion detection. Step 8 involves periodic activation of XAI. Finally, in Step 9, results of XAI and intrusion detection are sent to the CISO to enhance and debug the ML model for intrusion detection. Such a process is well-suited for the IoMT system since it continuously generates data, which enriches the ML used for IDS.
@@
3.4 XAI Porcess
ML has demonstrated its effectiveness in anomaly detection, particularly in identifying zero-day attacks and new vulnerabilities. This makes ML more advantageous compared to signature-based methods. However, non-transparent ML models, such as ANN, operate as black boxes, making it challenging to explain the reasoning behind the classification of instances, such as identifying an attack. This results in time consumption and ambiguity in analyzing predictions generated by these models. Explainable AI provides a solution by assisting model designers in specifying the impact of each attribute on the classification process. This proves beneficial for model designers, aiding in debugging, enhancing the ML model, gaining insights into its decision-making process, and fostering increased trust in the model.
@@ 3.4 XAI Porcess
4 EXPIREMENT SETUP AND RESULTS
@@
4.1 Hardwear Description
The experimentation was conducted utilizing Google Colab Pro as the designated testing environment. This cloud-based platform, constructed upon the foundation of Jupyter Notebook, offers collaborative access and augmented computational capabilities through GPU-equipped virtual machines. The specific environment, operating on the Linux operating system, was endowed with 12.7 GB of RAM and a storage capacity of 166.8 GB, facilitating the efficient management of data and computations.
To support the research endeavor, an array of essential packages was employed, encompassing TensorFlow, Pandas, NumPy, scikit-learn, SHAP for XAI, and Flower—a Framework for FL. The latter holds the advantage of executing on a large-scale FL [54] and can be employed to compare the outcomes of this study with those utilizing the same package. These packages assumed a pivotal role in the implementation and analysis of the experiments, furnishing a diverse range of tools and functionalities.

### v2

@@ 2.2 Related Work
In the realm of IDS-based ML incorporating XAI, a noteworthy instance is evident in the work presented in [40], the authors proposed the use of bidirectional SRU (Simple Recurrent Unit) with skip connections as a method for detecting attacks within IoMT networks. This approach effectively mitigates issues related to vanishing gradients, enhancing both training time and performance in attack detection. They compared the results achieved using two recurrent model variants, LSTM, and GRU, leveraging the ToN-IoT dataset [20]. Furthermore, they compared their findings with prior studies and discovered that their solution outperformed in terms of accuracy, precision, recall, and F-score. Additionally, they conducted an analysis of important features using an XAI technique called Local Interpretable Model-Agnostic Explanations (LIME).
Examining the reviewed solutions reveals a notable absence of an integrated approach that effectively leverages the advantages of ML for intrusion detection, FL for privacy protection, and XAI for interpretation and explanation within IoMT. Existing works focus on ML-based centralized IDS [36, 40, 14, 16, 21, 24, 26, 29, 30, 31, 35], where a central server collects and processes all data, leading to significant privacy risks and potential bottlenecks in data handling. While some studies have utilized DL techniques to improve the detection of cyber threats, these methods often lack transparency, operating as ”black-box” models that hinder trust and interpretability. Additionally, previous research on FL-based IDS solutions has primarily emphasized distributed learning efficiency but has failed to integrate mechanisms capable of explaining the decisions made by the models [37, 38, 39]. Although FL effectively addresses privacy concerns by decentralizing data processing, it does not inherently enhance the transparency of ML models. Furthermore, most reviewed solutions either exclude XAI methods entirely or rely exclusively on local model explanations [40]. In this context, the proposed solution in this article stands out as the initial effort to combine these diverse technologies ML, FL, and XAI into a cohesive framework tailored to enhance the security and interpretability of IoMT systems.
Ref Methods Datasets Learning approach use of the XAI Optimizing FL parameter
@@
[36] differential privacy contractive deep autoencoder for data fusion and quantum DNN method for intrusion detection WUSTL-EHMS [17] and ICU [19] central No No
[37] DNN CICIDS2017 [26] FL No No
[38] ANN MIMIC [38] FL No Yes
[39] hierarchical long-term memory ToN-IoT and NSL-KDD [15] FL No No
[40] bidirectional SRU with skip connections and LIME for XAI ToN-IoT [20] central Yes No
[14] PSO for features selections and AdaBoost for intrusion detection NSL-KDD [15] central No No
[16] RFE for features selections then MLP for intrusion detection WUSTL-EHMS [17], ECU-IoHT [18], ICU [19], and ToN-IoT [20] central No No
@@
[21] sSAE for dimensiality reduction then SGRU for intrusion detection AWID [22] central No No
@@
[24] multidimensional DL model for features selection based on CNN, CNN-LSTM, and bidirectional LSTM then fully connected layers for intrusion detection KISTI, KDDCup-99 [25], CICIDS2017 [26], WSN-DS [27] and UNSW-NB15 [28] central No No
[29] PSO for features selections then DNN for intrusion detection WUSTL-EHMS [17] central No No
@@
[30] integrates the LRGU technique into the MIFS. for features selection then ML for intrusion detection WUSTL-EHMS [17] central No No
@@
[31] PCA then GWO to minimize features then DNN for intrusion detection Kaggle intrusion data samples central No No
[32] ensemble learning that include NB, DT and RF ToN-IoT [20] Fog-Cloud No No
@@
[33] Deep Hierarchical Stacked Neural Networks UNSW-BOT-IoT [34] and UNSW-NB15 [28] Multi-Cloud No No
[35] ANN, RF, KNN, and SVM WUSTL-EHMS [17] central No No
@@ 3 METHODOLOGY FOR PROPOSED IDS MODEL
In this section, the IoMT system under consideration for the proposed framework is explored, followed by an outline of the FL process, which involves local training at the end-device level and aggregation at the server level. Subsequently, the employed XAI method is detailed. The objective of this framework is to ensure that data collection, transfer, and processing are conducted securely, respecting their privacy and aligning with international standards in the medical domain.
@@ 3.1 IoMT Systems
In layer 1, various types of medical devices equipped with sensors and actuators are employed to collect vital information and administer medications to patients that suffer from chronic diseases and also it can be used for fall detection for elderly persons or measuring the performance of athletes. These devices can be categorized into four types [41, 42]: implanted devices within the body, wearable devices, ambient devices capturing environmental data, and stationary devices found within hospitals. Given the energy limitations, wireless connections are established between these medical devices and mobile devices using low, or ultra-low-power wireless communication technologies such as Bluetooth, Zigbee, or NFC, thereby overcoming communication constraints [43].
@@
Finally, in layer 3, the centralized medical server is responsible for handling messages transmitted from the mobile devices and relaying them back to the patients. The server must be equipped with high computational capacity to effectively handle incoming and outgoing communications as well as perform in-depth analysis of the received data using AI methods. Additionally, it features a cloud server for intelligent decision-making, aggregating and storing additional patient medical data. The collected data is accessible to doctors, patients, and the pharmacy department through an online interface or smartphone. Integration with Electronic Health Record (EHR) and Electronic Medical Record (EMR) systems ensures easy access to information and provides notifications for uploaded or received health data. The figure 2 provides an overview of the described system [48].
@@ 3.2 Local Training Process
In this context, the inclination is towards employing supervised learning in DL for IDS, leveraging ANN for real-time anomaly detection. This architecture, simpler and less intricate than alternative models, enables adaptability and maintainability while delivering outstanding performance. Its capability to conduct real-time detection, as showcased in section 4, underscores its efficiency. Furthermore, the model’s straightforward design helps curb energy consumption within personal devices while contributing to reduced communication costs in the FL environment, which arise from exchanging model weights during the FL process instead of raw data. There are three types of neural layer in an ANN: an input layer, one or more hidden layers and an output layer. Each neuron contains a threshold and connections to other neurons with weighted connections. Neurons get activated if their cumulative weight surpasses the threshold, transmitting signals to the subsequent layer [50].
The rectified linear unit (ReLU) introduced in [51] serves as the activation function for the hidden layer, utilizing the MAX function (1) to enable faster computation, prevent overfitting, and enhance overall model performance, while the He-Initialization method from [52] is preferred for weight initialization due to its compatibility with ReLU (2), ensuring efficient training dynamics. For the output layer, the sigmoid activation function (3) is applied, making it well-suited for binary classification tasks. During model compilation, cross-entropy is employed as the loss function (4), and the Adam optimizer, described in [53] (5), is chosen for its efficiency, combining the benefits of adagrad, which handles sparse gradients effectively, and RMSProp, which excels in optimizing non-stationary objectives, while its low memory requirements make it particularly suitable for large-scale datasets or models with a high number of parameters [53].
@@
Where:
[…]
@@
Where:
[…]
@@
Where:
$g_{t}:\text{ is the gradient.}$.
$\alpha\ :\text{ is the learning rate.}$.
$\beta_{1}\beta_{2}:\text{ are the moment parameters.}$.
[…]
@@ 3.3 FL Porcess
Improving the security of the IoMT is crucial, and FL emerges as a pivotal solution. FL facilitates model training by allowing the exchange of locally trained model weights among end-devices, avoiding the transmission of raw data. This innovative approach upholds data privacy standards, aligning seamlessly with regulations such as HIPAA and GDPR, and its decentralized nature mitigates the vulnerability of a centralized structure.
@@
Input: Initial model weights $w_{0}$, number of communication rounds $R$, number of local epochs $E$, fraction of clients $Fr$, number of patients $M$, learning rate $\eta$.
Output: Updated global model weights $w_{t+1}$ after aggregation.
Description: The hospital server initializes the global model with random weights using He-Initialization. It then iterates through a series of communication rounds, where a random subset of clients (personal devices) is selected to participate in training. Each client updates its local model using its own data and sends the updated weights back to the server. The server aggregates these local updates by averaging them to produce a new global model. This process repeats until the model converges or reaches the desired performance.
Initialize model $w_{0}$ with He-Initialization $\triangleright$ Initialize global model weights using He-Initialization for efficient training.
$R\leftarrow$ Number of Round of communication $\triangleright$ Set the total number of communication rounds between the server and clients.
$E\leftarrow$ Number of local epoch $\triangleright$ Set the number of local training epochs for each client.
$Fr\leftarrow$ Fraction fit $\triangleright$ Set the fraction of clients to be selected in each communication round.
$M\leftarrow$ Number of Patient $\triangleright$ Set the total number of patients (clients) in the system.
$\eta\leftarrow$ learning rate $\triangleright$ Set the learning rate for model updates.
for $R=1,2,3\dots$ do $\triangleright$ Begin communication rounds.
     $C\leftarrow$ Random Set of $M\times Fr$ clients $\triangleright$ Randomly select a subset of clients to participate in the current round.
     for patient $k\subset C$ in parallel do $\triangleright$ Each selected client performs local training in parallel.
         $w_{(t+1)}^{k}\leftarrow$ Patient Client Update $(k,w_{t},E)$triangleright Client $k$ updates its local model using its data and sends the updated weights back to the server.
@@
     $w_{(t+1)}\leftarrow\sum_{k+1}^{k}\frac{\eta_{k}}{\eta}w_{(t+1)}^{k}$triangleright Aggregate the local model weights from all participating clients to update the global model.
@@
Algorithm 2 Patient Client Update
@@
Input: Current global model weights $w_{t}$, number of local epochs $E$, local data on the client device.
Output: Updated local model weights $w_{t+1}^{k}$ after local training.
Description: Each selected client performs local training on its data for $E$ epochs using the Adam optimizer to minimize the binary cross-entropy loss function. The client updates its local model weights based on the training data and sends the updated weights back to the hospital server for aggregation.
for $e=0$ to $e=E-1$ do $\triangleright$ Perform local training for $E$ epochs.
     $w\leftarrow$ use Adam optimizer to update $w$ to minimize the binary cross-entropy loss function $\triangleright$ Update local model weights using the Adam optimizer.
end for
return $w$ to server $\triangleright$ Send the updated local model weights back to the server.
@@
In Step 1, the security administrator initializes the global model with random weights at the hospital server. Personal devices selected for FL download the initial weights to initialize their local models. In Step 2, medical sensors send captured data to paired personal devices using short-range wireless communication. In Step 3, the local model undergoes training to update weights, utilizing a specified number of epochs as outlined in Algorithm 2, once personal devices have accumulated a sufficient amount of data. This strategy is designed to minimize the number of communication rounds needed for model convergence, thereby reducing communication costs through the optimization of bandwidth usage. In Step 4, the updated weights of local models are transmitted to the hospital server through encrypted communication protocols. Step 5 sees the hospital server aggregating these weights to update the global model by summing the received weights and dividing by the number of selected participants in a communication round, thereby creating the global model as depicated in formula (6). Step 6 involves the global server sending updated weights to personal devices, and Steps 2-6 are repeated until the weight modifications become insignificant or the model converges. Once the model achieves the desired performance in Step 7, it is deployed for intrusion detection. Step 8 involves periodic activation of XAI. Finally, In Step 9, the outcomes derived from XAI methodologies, along with the historical data pertaining to intrusion detection, are transmitted to the CISO either at predefined periodic intervals or upon request. This transmission facilitates the iterative refinement and debugging of the ML model employed for intrusion detection. The historical intrusion detection data is made accessible to users via their personnel devices, thereby augmenting their confidence in the efficacy and reliability of the detection model. Concurrently, regulatory bodies are granted access to both the XAI results and the intrusion detection history. This access ensures compliance with internationally recognized standards and verifies adherence to ethical principles, particularly within the healthcare sector.
Such a process is well-suited for the IoMT system since it continuously generates data, which enriches the ML models used for IDS.
@@
3.4 XAI Process
ML has demonstrated its effectiveness in anomaly detection, particularly in identifying zero-day attacks and new vulnerabilities. This makes ML more advantageous compared to signature-based methods. However, non-transparent ML models, such as ANN, operate as black boxes, making it challenging to explain the reasoning behind the classification of instances, such as identifying an attack. This results in time consumption and ambiguity in analyzing predictions generated by these models. Explainable AI provides a solution by assisting model designers in determining the impact of each attribute on the classification process. This proves beneficial for model designers, aiding in debugging, enhancing the ML model, gaining insights into its decision-making process, and fostering increased trust in the model.
@@ 3.4 XAI Process
3.5 Ethical Considerations
The proposed framework makes significant contributions to ethical considerations in healthcare by addressing core principles such as data protection, transparency, and accountability.
  • Data Protection and Privacy: The framework leverages FL for IDS, ensuring that patient data remains localized on their devices. Only model weights are securely shared, minimizing privacy risks and ensuring compliance with stringent regulations like HIPAA and GDPR. This approach reduces the likelihood of data breaches and cyberattacks, thereby fostering greater trust in the system.
  • Transparency in AI Decisions: By integrating XAI methods such as SHAP, the framework provides clear and interpretable explanations for the model’s decisions. This transparency enables stakeholders—including model designers, healthcare providers, and regulators—to understand how decisions are made, ensuring that AI outcomes are justifiable and aligned with user expectations and regulatory requirements.
  • Accountability and Bias Mitigation: The framework prioritizes fairness by utilizing diverse datasets and applying XAI techniques to detect and correct biases. This ensures that predictions are equitable and that any errors or biases in the model can be promptly identified and addressed. By doing so, it enhances accountability among system designers and operators, promoting responsible AI deployment.
  • Compliance with International Standards: The framework adheres to global standards such as ISO/IEC 27001 and ISO/IEC 27701, incorporating robust security measures like encrypted communication and decentralized data storage. Additionally, XAI ensures the traceability and transparency of model decisions, further reinforcing the framework’s credibility and reliability in a global context.
4 EXPERIMENT SETUP AND RESULTS
@@
The practical challenges in FL, such as device heterogeneity, network latency, and connectivity, are critical for real-world deployment. To streamline the analysis, several simplifying assumptions are introduced. First, it is presumed that the IoMT devices involved in the FL process have sufficient computational capabilities to conduct local model training. While these devices may vary in processing power, memory, and energy constraints, they are assumed to meet the minimum requirements necessary for training local models. This assumption shifts the focus toward optimizing the FL process, excluding extreme cases of severely resource-constrained devices.
Second, the network latency between IoMT devices and the central server is assumed to remain within acceptable thresholds for real-time communication. This ensures timely interactions during the FL process.
Finally, stable connectivity is presumed throughout the FL training phase. Although real-world scenarios may experience occasional disruptions, the devices are assumed to stay connected long enough to complete local training and exchange updates with the server. This allows the analysis to concentrate on the FL process itself, without addressing frequent disconnections or network instability.
4.1 Hardware Description
The experimentation is conducted utilizing Google Colab Pro as the designated testing environment. This cloud-based platform, constructed upon the foundation of Jupyter Notebook, offers collaborative access and augments computational capabilities through GPU-equipped virtual machines. The specific environment, operating on the Linux operating system, is endowed with 12.7 GB of RAM and a storage capacity of 166.8 GB, facilitating the efficient management of data and computations.
To support the research endeavor, an array of essential packages is employed, encompassing TensorFlow, Pandas, NumPy, scikit-learn, SHAP for XAI, and Flower—a Framework for FL. The latter holds the advantage of executing on a large-scale FL [56] and can be employed to compare the outcomes of this study with those utilizing the same package. These packages assume a pivotal role in the implementation and analysis of the experiments, furnishing a diverse range of tools and functionalities.